"Caspian Sea Drinks is considering the production of a diet drink. The expansion of the plant and the purchase of the equipment necessary to produce the diet drink will cost $23.00 million. The plant and equipment will be depreciated over 10 years to a book value of $3.00 million, and sold for that amount in year 10. Net working capital will increase by $1.43 million at the beginning of the project and will be recovered at the end. The new diet drink will produce revenues of $8.88 million per year and cost $2.20 million per year over the 10-year life of the project. Marketing estimates 12.00% of the buyers of the diet drink will be people who will switch from the regular drink. The marginal tax rate is 32.00%. The WACC is 13.00%. Find the NPV (net present value)."

2 decimal places
Answer #1

Profit = (revenues-sales)*(1-switch%)

=(8880000-2200000)*(1-0.12) = 5878400

Time line 0 1 2 3 4 5 6 7 8 9 10
Cost of new machine -23000000
Initial working capital -1430000
=Initial Investment outlay -24430000
100.00%
Profits 5878400 5878400 5878400 5878400 5878400 5878400 5878400 5878400 5878400 5878400
-Depreciation (Cost of equipment-salvage value)/no. of years -2000000 -2000000 -2000000 -2000000 -2000000 -2000000 -2000000 -2000000 -2000000 -2000000 3000000 =Salvage Value
=Pretax cash flows 3878400 3878400 3878400 3878400 3878400 3878400 3878400 3878400 3878400 3878400
-taxes =(Pretax cash flows)*(1-tax) 2637312 2637312 2637312 2637312 2637312 2637312 2637312 2637312 2637312 2637312
+Depreciation 2000000 2000000 2000000 2000000 2000000 2000000 2000000 2000000 2000000 2000000
=after tax operating cash flow 4637312 4637312 4637312 4637312 4637312 4637312 4637312 4637312 4637312 4637312
reversal of working capital 1430000
+Proceeds from sale of equipment after tax =selling price* ( 1 -tax rate) 2040000
+Tax shield on salvage book value =Salvage value * tax rate 960000
=Terminal year after tax cash flows 4430000
Total Cash flow for the period -24430000 4637312 4637312 4637312 4637312 4637312 4637312 4637312 4637312 4637312 9067312
Discount factor= (1+discount rate)^corresponding period 1 1.13 1.2769 1.442897 1.6304736 1.8424352 2.0819518 2.35260548 2.6584442 3.004042 3.394567
Discounted CF= Cashflow/discount factor -24430000 4103815.929 3631695.513 3213889.834 2844150.3 2516947.2 2227386.9 1971138.824 1744370.6 1543691 2671124
NPV= Sum of discounted CF= 2038210.37